# Cluster algebras and quantum cohomology rings: A-type

### v1

@@ 5.3 Quantum cohomology ring of $Fl(N_{1},\ldots,N_{n+1})$
$\prod_{i=1}^{N_{k}}(t-x^{(k)}_{i})\sum_{b=0}^{N_{l}-N_{k}}e_{b}(-\vec{x}^{(l)})\left[\sum_{a=0}^{N_{l}-N_{k}-b}t^{a}h_{N_{l}-N_{k}-a-b}(\vec{x}^{(k)})\right]\omega\,.$ (5.30)
The formula inside the bracket is actually $h_{N_{l}-N_{k}-b}(t,\vec{x}^{(k)})$. We view $t$ as $x_{0}^{(k)}$, and then $\prod_{i=1}^{N_{k}}(t-x^{(k)}_{i})\omega_{k}$ is the Vandermonde formula of variables $(x^{(k)}_{i})_{i=0}^{N_{k}}$. The formula (5.30) can then be written as follows,
$\sum_{b=0}^{N_{l}-N_{k}}e_{b}(-\vec{x}^{(l)})\det\begin{bmatrix}t^{N_{l}-b}&\left(x_{1}^{(k)}\right)^{N_{l}-b}&\cdots&\left(x_{N_{k}}^{(k)}\right)^{N_{l}-b}\\ t^{N_{k}-1}&\left(x_{1}^{(k)}\right)^{N_{k}-1}&\cdots&\left(x_{N_{k}}^{(k)}\right)^{N_{k}-1}\\ \vdots&\vdots&\ddots&\vdots\\ 1&1&\cdots&1\\ \end{bmatrix}\prod_{m\neq k}\omega_{m}$ (5.31)
Notice that the determinant is zero once $b$ is greater than $N_{l}-N_{k}$, so we can change the summation index from 00 to $N_{l}$. Multiply $e_{b}(-\vec{x}^{(l)})$ and $\prod_{m\neq k}\omega_{m}$ to the first row of the determinant. Then entries in the 1st row and $j$-th column for $j\geq 2$ are $\prod_{i=1}^{N_{l}}(x^{(k)}_{j-1}-x^{(l)}_{i})\omega_{m\neq k}$ which are all zero by Lemma 5.7. The formula (5.31) becomes $\prod_{i=1}^{N_{l}}(t-x^{(l)}_{i})\omega=\widetilde{c_{t}^{S}(S_{l})}\cup\omega$. ∎
For our purpose, we rewrite the lifting $\widetilde{c_{t}^{S}(S_{l}/S_{k})}\omega$ as a linear combination of elements in (5.14). For $\vec{x}=(x_{1},x_{2},\ldots,x_{n})$ and $\vec{m}=(m_{1},m_{2},\ldots,m_{n})\in\mathbb{Z}^{n}$, denote
$D^{\vec{m}}(\vec{x}):=\det\begin{bmatrix}x_{1}^{m_{n}}&x_{2}^{m_{n}}&\cdots&x_{n}^{m_{n}}\\ x_{1}^{m_{n-1}}&x_{2}^{m_{n-1}}&\cdots&x_{n}^{m_{n-1}}\\ \vdots&\vdots&\ddots&\vdots\\ x_{1}^{m_{1}}&x_{2}^{m_{1}}&\cdots&x_{n}^{m_{1}}\\ \end{bmatrix}\,.$ (5.32)
@@
$\widetilde{c_{t}^{S}(S_{l}/S_{k})}\omega=\sum_{a=0}^{N_{l}-N_{k}}t^{a}\left[\sum_{b=0}^{N_{l}-N_{k}-a}e_{b}(-\vec{x}^{(l)})\sum_{0\leq m_{i}<N_{i+1};i=k,\ldots,l-1}(-1)^{\bullet}D^{\vec{m}_{k}}(\vec{x}^{(k)})D^{\vec{m}_{k+1}}(\vec{x}^{(k+1)})\cdots D^{\vec{m}_{l}}(\vec{x}^{(l)})\right]\prod_{j<k\text{ or }j>l}\omega_{j}.$ (5.33)
@@
By (5.28), we have
$\widetilde{c_{t}^{S}(S_{l}/S_{k})}\omega=\sum_{a=0}^{N_{l}-N_{k}}t^{a}\left[\sum_{b=0}^{N_{l}-N_{k}-a}e_{b}(-\vec{x}^{(l)})D^{(0,1,\ldots,N_{k}-2,N_{l}-a-b-1)}(\vec{x}^{(k)})\right]\prod_{j\neq k}\omega_{j}.$ (5.34)
@@
Notice that the left hand side of the above equation vanishes by (5.12), so does the right hand side. Then we have
$\left(x_{i}^{(k)}\right)^{N_{l}-a-b-1}\omega_{k+1}=\sum_{m_{k}=0}^{N_{k+1}-1}(-1)^{N_{k+1}-m_{k}-1}\left(x_{i}^{(k)}\right)^{m_{k}}\cdot D^{(0,1,\ldots,m_{k}-1,m_{k}+1,\ldots,N_{k+1}-1,N_{l}-a-b-1)}(\vec{x}^{(k+1)})\,.$ (5.35)
@@
$D^{(0,1,\ldots,N_{k}-2,N_{l}-a-b-1)}(\vec{x}^{(k)})\prod_{j\neq k}\omega_{j}=\sum_{0\leq m_{i}<N_{i+1};\atop i=k,\ldots,l-1}(-1)^{\bullet}D^{\vec{m}_{k}}(\vec{x}^{(k)})D^{\vec{m}_{k+1}}(\vec{x}^{(k+1)})\cdots D^{\vec{m}_{l}}(\vec{x}^{(l)})\prod_{j<k\text{ or }j>l}\omega_{j}.$ (5.36)
Finally we only need to substitute (5.36) into (5.34). ∎
Although, the right hand side of (5.33) looks complicated, one can check that it is a linear combination of elements in the set (5.14). In particular, the degree of $x^{(i)}_{j}$ is less than $N_{i+1}$ for any $i,j$.
5.4 A fundamental relation in quantum cohomology of $Fl$
Recall that $Fl$ is a tower of Grassmannian $Gr(N_{1},N_{2})\rightarrow Gr(N_{2},N_{3})\rightarrow\cdots\rightarrow Gr(N_{n},N_{n+1})$ (cf. (4.16)). Let $[\ell_{k}]$ denote the rational curve class lying inside the $k$-th tower $Gr(N_{k},N_{k+1})$. Then the Novikov ring of $Fl$ is $N(Fl)=\bigoplus_{k}\mathbb{Z}_{\geq 0}[\ell_{k}]$. Set the Novikov variable $q_{k}:=q^{[\ell_{k}]}$. By [GK24] the specialization of Novikov variable in Proposition 5.5 is
$q_{k,i}\mapsto(-1)^{N_{k}-1}q_{k},$ (5.37)
@@ 5.4 A fundamental relation in quantum cohomology of $Fl$
We only have to prove that the quantum correction of the quantum product is $(-1)^{N_{k}+N_{k-1}}q_{k}c^{S}_{t}(S_{l}/S_{k+1})*c^{S}_{t}(S_{k-1})$ by utilizing Proposition 5.5. For $\gamma=c^{S}_{t}(S_{k})$ and the lifting $\widetilde{\gamma}$ in (5.21), we need to find $\zeta\in H_{S}^{*}(V\sslash_{\theta}T)^{W}$ such that $\zeta*\omega=\widetilde{\gamma}\cup\omega$. Actually, we have $\widetilde{\gamma}*\omega=\widetilde{\gamma}\cup\omega$ by degree reason. Therefore,
$\zeta=\widetilde{\gamma}=\prod_{j=1}^{N_{k}}(t-x^{(k)}_{j})=\sum_{l=0}^{N_{k}}t^{N_{k}-l}e_{l}(-\vec{x}^{(k)}).$ (5.39)
For $\gamma^{\prime}=c^{S}_{t}(S_{l}/S_{k})$, by (5.33) we have
$\widetilde{\gamma^{\prime}}\cup\omega=\sum_{a=0}^{N_{l}-N_{k}}t^{a}\left[\sum_{b=0}^{N_{l}-N_{k}-a}e_{b}(-\vec{x}^{(l)})\sum_{0\leq m_{i}<N_{i+1};\atop i=k,\ldots,l-1}(-1)^{\bullet}D^{\vec{m}_{k}}(\vec{x}^{(k)})D^{\vec{m}_{k+1}}(\vec{x}^{(k+1)})\cdots D^{\vec{m}_{l}}(\vec{x}^{(l)})\right]\prod_{j<k\text{ or }j>l}\omega_{j}.$ (5.40)
The quantum correction $\xi*(\widetilde{\gamma^{\prime}}\cup\omega)$ in $QH^{*}(V\sslash_{\theta}T)$ under the specialization (5.37) can be computed as follows. The quantum correction only happens when $m_{k}=N_{k+1}-1$,
@@
$\left.\sum_{0\leq m_{i}<N_{i+1};\atop i=k+1,\ldots,l-1}(-1)^{\bullet}D^{\vec{m}_{k+1}}(\vec{x}^{(k+1)})\cdots D^{\vec{m}_{l}}(\vec{x}^{(l)})\right)\prod_{j<k\text{ or }j>l}\omega_{j}.$ (5.41)
@@
$\sum_{l=0}^{N_{k}}t^{N_{k}-l}e_{l}(-\vec{x}^{(k)})*\begin{bmatrix}\left(x_{1}^{(k)}\right)^{N_{k+1}-1}&\cdots&\left(x_{N_{k}}^{(k)}\right)^{N_{k+1}-1}\\ \left(x_{1}^{(k)}\right)^{N_{k}-2}&\cdots&\left(x_{N_{k}}^{(k)}\right)^{N_{k}-2}\\ \vdots&\ddots&\vdots\\ 1&\cdots&1\\ \end{bmatrix}$ (5.42)
@@
$\sum_{l=0}^{N_{k}}t^{N_{k}-l}e_{l}(-\vec{x}^{(k)})*\sum_{i=0}^{N_{k+1}-N_{k}}t^{i}\begin{bmatrix}\left(x_{1}^{(k)}\right)^{N_{k+1}-1-i}&\cdots&\left(x_{N_{k}}^{(k)}\right)^{N_{k+1}-1-i}\\ \left(x_{1}^{(k)}\right)^{N_{k}-2}&\cdots&\left(x_{N_{k}}^{(k)}\right)^{N_{k}-2}\\ \vdots&\ddots&\vdots\\ 1&\cdots&1\\ \end{bmatrix}\,.$ (5.43)
The formula in (5.43) is equal to
@@
$-\sum_{i=1}^{N_{k}}\left(\left(x_{i}^{(k)}\right)^{*N_{k+1}}-\left(x_{i}^{(k)}\right)^{N_{k+1}}\right)\det A_{i}(t)=\sum_{i=1}^{N_{k}}(-1)^{N_{k}}q_{k}\prod_{a=1}^{N_{k-1}}(x^{(k-1)}_{a}-x^{(k)}_{i})\det A_{i}(t)$ (5.45)
@@
The formula (5.45) and hence the quantum correction of (5.42) by Lagrangian interpolation is $(-1)^{N_{k}+N_{k-1}}q_{k}\widetilde{c^{S}_{t}(S_{k-1})}\omega_{k}$. The remaining part of (5.4) excluding (5.42) is $\widetilde{c^{S}_{t}(S_{l}/S_{k+1})}(\omega/\omega_{k})$. Then one gets the quantum correction of (5.4) is $(-1)^{N_{k}+N_{k-1}}q_{k}\widetilde{c^{S}_{t}(S_{k-1})}*\widetilde{c^{S}_{t}(S_{l}/S_{k+1})}\omega$.
The reduction of (5.42) via (5.6) will also produce a classical part. A quantum correction may appear in the quantum multiplication between this classical part and the remaining part of (5.4). We claim this quantum correction vanishes. In fact, the classical part is a summation of the following polynomials,
$f(t,x^{(k)}_{1},\ldots,x^{(k)}_{i-1},x^{(k)}_{i-1},x^{(k)}_{i-1},\ldots,x^{(k)}_{N_{k}})\cdot\sum_{l=1}^{N_{k}}\left(x_{i}^{(k)}\right)^{N_{k}-l}e_{l}(-\vec{x}^{(k+1)}),\quad i=1,2,\ldots,N_{k}$ (5.46)
@@
$\sum_{l=1}^{N_{k}}\left(x_{i}^{(k)}\right)^{N_{k}-l}e_{l}(-\vec{x}^{(k+1)})*\begin{bmatrix}\left(x_{1}^{(k+1)}\right)^{N_{k+2}-1}&\cdots&\left(x_{N_{k}}^{(k+1)}\right)^{N_{k+2}-1}\\ \left(x_{1}^{(k+1)}\right)^{N_{k+1}-2}&\cdots&\left(x_{N_{k}}^{(k+1)}\right)^{N_{k+1}-2}\\ \vdots&\ddots&\vdots\\ 1&\cdots&1\\ \end{bmatrix}$ (5.47)
Via the same procedure as in (5.42) (just by replacing $t$ by $x_{i}^{(k)}$), one can prove that the quantum correction has a factor $(-1)^{\bullet}q_{k+1}\prod_{l=1}^{N_{k}}(x_{i}^{(k)}-x^{(k)}_{l})=0$. (One may be concerned about the summation index in (5.47) starting from $l=1$, which is different from (5.42). In fact, one can change $l=0$ to $l=1$ in (5.42) without changing the quantum correction). Hence all quantum corrections in the reduction procedure vanish except the first one. Therefore the quantum correction of $c^{S}_{t}(S_{k})*c^{S}_{t}(S_{l}/S_{k})$ is $(-1)^{N_{k}+N_{k-1}}q_{k}c^{S}_{t}(S_{k-1})*c^{S}_{t}(S_{l}/S_{k+1})$. By (5.23), the classical part is $c^{S}_{t}(S_{l})$, which completes the proof. ∎
@@
In the special case $l=k+1$, the formula (5.38) originally appears in a physical literature [Wit93] for $Fl=Gr(k,N)$, and it is also discussed in [GMS+24, Section 4.6] for general partial flag varieties.
@@
According to (5.38), the left hand side is equal to
@@
which is equal to the right hand side. ∎

### v2

@@ 5.3 Quantum cohomology rings of flag varieties
(51) $\prod_{i=1}^{N_{k}}(t-x^{(k)}_{i})\sum_{b=0}^{N_{l}-N_{k}}e_{b}(-\vec{x}^{(l)})\left[\sum_{a=0}^{N_{l}-N_{k}-b}t^{a}h_{N_{l}-N_{k}-a-b}(\vec{x}^{(k)})\right]\omega\,.$
The formula inside the bracket is actually $h_{N_{l}-N_{k}-b}(t,\vec{x}^{(k)})$. We view $t$ as $x_{0}^{(k)}$, and then $\prod_{i=1}^{N_{k}}(t-x^{(k)}_{i})\omega_{k}$ is the Vandermonde formula of variables $(x^{(k)}_{i})_{i=0}^{N_{k}}$. The formula (51) can then be written as follows,
(52) $\sum_{b=0}^{N_{l}-N_{k}}e_{b}(-\vec{x}^{(l)})\det\begin{bmatrix}t^{N_{l}-b}&\left(x_{1}^{(k)}\right)^{N_{l}-b}&\cdots&\left(x_{N_{k}}^{(k)}\right)^{N_{l}-b}\\ t^{N_{k}-1}&\left(x_{1}^{(k)}\right)^{N_{k}-1}&\cdots&\left(x_{N_{k}}^{(k)}\right)^{N_{k}-1}\\ \vdots&\vdots&\ddots&\vdots\\ 1&1&\cdots&1\\ \end{bmatrix}\prod_{m\neq k}\omega_{m}$
Notice that the determinant is zero once $b$ is greater than $N_{l}-N_{k}$, so we can change the summation index from 00 to $N_{l}$. Multiply $e_{b}(-\vec{x}^{(l)})$ and $\prod_{m\neq k}\omega_{m}$ to the first row of the determinant. Then entries in the 1st row and $j$-th column for $j\geq 2$ are $\prod_{i=1}^{N_{l}}(x^{(k)}_{j-1}-x^{(l)}_{i})\omega_{m\neq k}$ which are all zero by Lemma 5.7. The formula (52) becomes $\prod_{i=1}^{N_{l}}(t-x^{(l)}_{i})\omega=\widetilde{c_{t}^{S}(S_{l})}\cup\omega$. ∎
For our purpose, we rewrite the lifting $\widetilde{c_{t}^{S}(S_{l}/S_{k})}\omega$ as a linear combination of elements in (42). For $\vec{x}=(x_{1},x_{2},\ldots,x_{n})$ and $\vec{m}=(m_{1},m_{2},\ldots,m_{n})\in\mathbb{Z}^{n}$, denote
$D^{\vec{m}}(\vec{x}):=\det\begin{bmatrix}x_{1}^{m_{n}}&x_{2}^{m_{n}}&\cdots&x_{n}^{m_{n}}\\ x_{1}^{m_{n-1}}&x_{2}^{m_{n-1}}&\cdots&x_{n}^{m_{n-1}}\\ \vdots&\vdots&\ddots&\vdots\\ x_{1}^{m_{1}}&x_{2}^{m_{1}}&\cdots&x_{n}^{m_{1}}\\ \end{bmatrix}\,.$
@@
(53) $\widetilde{c_{t}^{S}(S_{l}/S_{k})}\omega=\sum_{a=0}^{N_{l}-N_{k}}t^{a}\left[\sum_{b=0}^{N_{l}-N_{k}-a}e_{b}(-\vec{x}^{(l)})\sum_{0\leq m_{i}<N_{i+1};i=k,\ldots,l-1}(-1)^{\bullet}D^{\vec{m}_{k}}(\vec{x}^{(k)})D^{\vec{m}_{k+1}}(\vec{x}^{(k+1)})\cdots D^{\vec{m}_{l}}(\vec{x}^{(l)})\right]\prod_{j<k\text{ or }j>l}\omega_{j}.$
@@
By (50), we have
(54) $\widetilde{c_{t}^{S}(S_{l}/S_{k})}\omega=\sum_{a=0}^{N_{l}-N_{k}}t^{a}\left[\sum_{b=0}^{N_{l}-N_{k}-a}e_{b}(-\vec{x}^{(l)})D^{(0,1,\ldots,N_{k}-2,N_{l}-a-b-1)}(\vec{x}^{(k)})\right]\prod_{j\neq k}\omega_{j}.$
@@
Notice that the left-hand side of the above equation vanishes by (41), so does the right-hand side. Then we have
$\left(x_{i}^{(k)}\right)^{N_{l}-a-b-1}\omega_{k+1}=\sum_{m_{k}=0}^{N_{k+1}-1}(-1)^{N_{k+1}-m_{k}-1}\left(x_{i}^{(k)}\right)^{m_{k}}\cdot D^{(0,1,\ldots,m_{k}-1,m_{k}+1,\ldots,N_{k+1}-1,N_{l}-a-b-1)}(\vec{x}^{(k+1)})\,.$
@@
(55) $D^{(0,1,\ldots,N_{k}-2,N_{l}-a-b-1)}(\vec{x}^{(k)})\prod_{j\neq k}\omega_{j}=\sum_{\begin{subarray}{c}0\leq m_{i}<N_{i+1};\\ i=k,\ldots,l-1\end{subarray}}(-1)^{\bullet}D^{\vec{m}_{k}}(\vec{x}^{(k)})D^{\vec{m}_{k+1}}(\vec{x}^{(k+1)})\cdots D^{\vec{m}_{l}}(\vec{x}^{(l)})\prod_{j<k\text{ or }j>l}\omega_{j}.$
Finally, we only need to substitute (55) into (54). ∎
Although, the right-hand side of (53) looks complicated, one can check that it is a linear combination of elements in the set (42). In particular, the degree of $x^{(i)}_{j}$ is less than $N_{i+1}$ for any $i,j$.
5.4 Fundamental relations in quantum cohomology rings of flag varieties
Recall that $Fl$ is a tower of Grassmannian $Gr(N_{1},N_{2})\rightarrow Gr(N_{2},N_{3})\rightarrow\cdots\rightarrow Gr(N_{n},N_{n+1})$ (cf. (22)). Let $[\ell_{k}]$ denote the rational curve class lying inside the $k$-th tower $Gr(N_{k},N_{k+1})$. Then the Novikov ring of $Fl$ is $N(Fl)=\bigoplus_{k}\mathbb{Z}_{\geq 0}[\ell_{k}]$. Set the Novikov variable $q_{k}:=q^{[\ell_{k}]}$. By [rimhookrule] the specialization of Novikov variable in Proposition 5.5 is
(56) $q_{k,i}\mapsto(-1)^{N_{k}-1}q_{k},$
@@ 5.4 Fundamental relations in quantum cohomology rings of flag varieties
We only have to prove that the quantum correction of the quantum product is $(-1)^{N_{k}+N_{k-1}}q_{k}c^{S}_{t}(S_{l}/S_{k+1})*c^{S}_{t}(S_{k-1})$ by utilizing Proposition 5.5. Z For $\gamma=c^{S}_{t}(S_{k})$ and the lifting $\widetilde{\gamma}$ in (46), we need to find $\zeta\in H_{S}^{*}(V\sslash_{\theta}T)^{W}$ such that $\zeta*\omega=\widetilde{\gamma}\cup\omega$. Actually, we have $\widetilde{\gamma}*\omega=\widetilde{\gamma}\cup\omega$ by degree reason. Therefore,
$\zeta=\widetilde{\gamma}=\prod_{j=1}^{N_{k}}(t-x^{(k)}_{j})=\sum_{l=0}^{N_{k}}t^{N_{k}-l}e_{l}(-\vec{x}^{(k)}).$
For $\gamma^{\prime}=c^{S}_{t}(S_{l}/S_{k})$, by (53) we have
$\widetilde{\gamma^{\prime}}\cup\omega=\sum_{a=0}^{N_{l}-N_{k}}t^{a}\left[\sum_{b=0}^{N_{l}-N_{k}-a}e_{b}(-\vec{x}^{(l)})\sum_{\begin{subarray}{c}0\leq m_{i}<N_{i+1};\\ i=k,\ldots,l-1\end{subarray}}(-1)^{\bullet}D^{\vec{m}_{k}}(\vec{x}^{(k)})D^{\vec{m}_{k+1}}(\vec{x}^{(k+1)})\cdots D^{\vec{m}_{l}}(\vec{x}^{(l)})\right]\prod_{j<k\text{ or }j>l}\omega_{j}.$
The quantum correction $\xi*(\widetilde{\gamma^{\prime}}\cup\omega)$ in $QH^{*}(V\sslash_{\theta}T)$ under the specialization (56) can be computed as follows. The quantum correction only happens when $m_{k}=N_{k+1}-1$,
@@
(58) $\left.\sum_{\begin{subarray}{c}0\leq m_{i}<N_{i+1};\\ i=k+1,\ldots,l-1\end{subarray}}(-1)^{\bullet}D^{\vec{m}_{k+1}}(\vec{x}^{(k+1)})\cdots D^{\vec{m}_{l}}(\vec{x}^{(l)})\right)\prod_{j<k\text{ or }j>l}\omega_{j}.$
@@
(59) $\sum_{l=0}^{N_{k}}t^{N_{k}-l}e_{l}(-\vec{x}^{(k)})*\det\begin{bmatrix}\left(x_{1}^{(k)}\right)^{N_{k+1}-1}&\cdots&\left(x_{N_{k}}^{(k)}\right)^{N_{k+1}-1}\\ \left(x_{1}^{(k)}\right)^{N_{k}-2}&\cdots&\left(x_{N_{k}}^{(k)}\right)^{N_{k}-2}\\ \vdots&\ddots&\vdots\\ 1&\cdots&1\\ \end{bmatrix}$
@@
(60) $\sum_{l=0}^{N_{k}}t^{N_{k}-l}e_{l}(-\vec{x}^{(k)})*\sum_{i=0}^{N_{k+1}-N_{k}}t^{i}\det\begin{bmatrix}\left(x_{1}^{(k)}\right)^{N_{k+1}-1-i}&\cdots&\left(x_{N_{k}}^{(k)}\right)^{N_{k+1}-1-i}\\ \left(x_{1}^{(k)}\right)^{N_{k}-2}&\cdots&\left(x_{N_{k}}^{(k)}\right)^{N_{k}-2}\\ \vdots&\ddots&\vdots\\ 1&\cdots&1\\ \end{bmatrix}\,.$
The formula in (60) is equal to
@@
(62) $-\sum_{i=1}^{N_{k}}\left(\left(x_{i}^{(k)}\right)^{*N_{k+1}}-\left(x_{i}^{(k)}\right)^{N_{k+1}}\right)\det A_{i}(t)=\sum_{i=1}^{N_{k}}(-1)^{N_{k}}q_{k}\prod_{a=1}^{N_{k-1}}(x^{(k-1)}_{a}-x^{(k)}_{i})\det A_{i}(t)$
@@
The formula (62) and hence the quantum correction of (59) by Lagrangian interpolation is $(-1)^{N_{k}+N_{k-1}}q_{k}\widetilde{c^{S}_{t}(S_{k-1})}\omega_{k}$. The remaining part of (5.4) excluding (59) is $\widetilde{c^{S}_{t}(S_{l}/S_{k+1})}(\omega/\omega_{k})$. Then one gets the quantum correction of (5.4) is $(-1)^{N_{k}+N_{k-1}}q_{k}\widetilde{c^{S}_{t}(S_{k-1})}*\widetilde{c^{S}_{t}(S_{l}/S_{k+1})}\omega$.
The reduction of (59) via (5.6) will also produce a classical part. A quantum correction may appear in the quantum multiplication between this classical part and the remaining part of (5.4). We claim this quantum correction vanishes. In fact, the classical part is a summation of the following polynomials,
$f(t,x^{(k)}_{1},\ldots,x^{(k)}_{i-1},x^{(k)}_{i-1},x^{(k)}_{i-1},\ldots,x^{(k)}_{N_{k}})\cdot\sum_{l=1}^{N_{k}}\left(x_{i}^{(k)}\right)^{N_{k}-l}e_{l}(-\vec{x}^{(k+1)}),\quad i=1,2,\ldots,N_{k}$
@@
(63) $\sum_{l=1}^{N_{k+1}}\left(x_{i}^{(k)}\right)^{N_{k+1}-l}e_{l}(-\vec{x}^{(k+1)})*\begin{bmatrix}\left(x_{1}^{(k+1)}\right)^{N_{k+2}-1}&\cdots&\left(x_{N_{k}}^{(k+1)}\right)^{N_{k+2}-1}\\ \left(x_{1}^{(k+1)}\right)^{N_{k+1}-2}&\cdots&\left(x_{N_{k}}^{(k+1)}\right)^{N_{k+1}-2}\\ \vdots&\ddots&\vdots\\ 1&\cdots&1\\ \end{bmatrix}$
Via the same procedure as in (59) (just by replacing $t$ by $x_{i}^{(k)}$), one can prove that the quantum correction has a factor $(-1)^{\bullet}q_{k+1}\prod_{l=1}^{N_{k}}(x_{i}^{(k)}-x^{(k)}_{l})=0$. (One may be concerned about the summation index in (63) starting from $l=1$, which is different from (59). In fact, one can change $l=0$ to $l=1$ in (59) without changing the quantum correction). Hence all quantum corrections in the reduction procedure vanish except the first one. Therefore the quantum correction of $c^{S}_{t}(S_{k})*c^{S}_{t}(S_{l}/S_{k})$ is $(-1)^{N_{k}+N_{k-1}}q_{k}c^{S}_{t}(S_{k-1})*c^{S}_{t}(S_{l}/S_{k+1})$. By (47), the classical part is $c^{S}_{t}(S_{l})$, which completes the proof. ∎
@@
In the special case $l=k+1$, the formula (57) originally appears in a physical literature [Witten] for $Fl=Gr(k,N)$, and it is also discussed in [MR4701783, Section 4.6] for general partial flag varieties.
@@
According to (57), the left-hand side is equal to
@@
which is equal to the right-hand side. ∎